# Shaken, Not Stirred: A Novel Dataset for Visual Understanding of Glasses in Human-Robot Bartending Tasks

### v2

@@ Abstract
Datasets for object detection often do not account for enough variety of glasses, due to their transparent and reflective properties. Specifically, open-vocabulary object detectors, widely used in embodied robotic agents, fail to distinguish subclasses of glasses. This scientific gap poses an issue to robotic applications that suffer from accumulating errors between detection, planning, and action execution. The paper introduces a novel method for the acquisition of real-world data from RGB-D sensors that minimizes human effort. We propose an auto-labeling pipeline that generates labels for all the acquired frames based on the depth measurements. We provide a novel real-world glass object dataset [^3] https://gajdosech2.github.io/GlassNICOLDataset/ that was collected on the Neuro-Inspired COLlaborator (NICOL), a humanoid robot platform. The data set consists of 7850 images recorded from five different cameras. We show that our trained baseline model outperforms state-of-the-art open-vocabulary approaches. In addition, we deploy our baseline model in an embodied agent approach to the NICOL platform, on which it achieves a success rate of 81% in a human-robot bartending scenario.
@@ I INTRODUCTION
Transparent objects are everywhere, from households, healthcare assistance, and gastronomy to industrial and construction-site environments. Yet, these materials reflect only a very small fraction of visible light. Most of it is scattered and passes through. From a computer vision perspective, the processing of these objects is challenging but crucial for a successful deployment in real-world robotics applications [1]. As the intelligent robot starts to make its way into public spaces, where glass objects in the form of bottles and drinking glasses are an integral item, future generations of service robots are required to not only detect but also manipulate those materials. In addition, machine learning and data-driven methods require a large amount of training data to reach a suitable performance. Similar to robotic Sim2Real applications, existing research also proves the presence of a significant Sim2Real gap in the area of synthetic glass material generation [2, 3, 1], while those approaches require sophisticated photorealistic rendering methods.
Open‑vocabulary approaches are used for object detection in a plethora of embodied robotic setups, such as [4]. Embodied robotic agents, at a minimum, consist of a detection module, an action module, and an LLM utilized for high-level interaction and task planning. Recent approaches aim to fuse detection and language processing by utilizing vision-language models (VLMs) [5]. Being interactive by design, these agents are typically situated in scenarios that involve user interaction and often also collaboration, thus leaving a potentially infinite copora of object names to be detected.
Well-known open-vocabulary object detection approaches such as Grounding DINO [6], YOLOWorld [7], and OWL-Vit [8] have revolutionized object detection by expanding the range of recognizable object classes without extensive re‑annotation. These approaches are built upon end-to-end transformer‑based architectures. Global image features are extracted by decoder backbones and fused with language embeddings by utilizing deep metric learning. However, despite their success, our experiments suggest that these models exhibit a significant decrease in performance when encountering transparent objects, commonly found in HRI scenarios. To address these challenges, we introduce a novel real-world dataset that was captured on the humanoid robot NICOL shown in Fig. 1, including a new depth-based automated labeling method. The robot is equipped with a combination of RGB-D scanners and standard RGB cameras. Specific contributions of our work include:
@@ II-A Transparent Material Perception
Missing or incorrect RGB-D data for robotic perception can be estimated using implicit functions and geometric priors [9] and by balancing local and global depth features [10]. The depth map can also be de-projected into a point cloud and processed with techniques like 3D CNN to achieve depth completion [11]. Iterative refinement is also possible, with attention module on transparent regions [12] or with indirect geometry representations [13].
RGB often contains information missing from depth data, which can be leveraged with transfer learning [14]. Cues can be extracted from the RGB to fill-in depth observations using global optimization [1]. Incorporating affordance detection of the underlying object yield further improvement [15]. Other approaches include multimodality, such as visual-tactile fusion [16] and usage of polarization imaging [17].
Recently, neural radiance fields (NeRFs), have emerged as a promising tool for transparent object perception. GraspNeRF introduces a multiview-based 6-DoF grasp detection system [18]. Evo-NeRF extends this approach with real-time NeRF training and grasp adaptation [19]. Multi-view methods enhance transparent object perception by integrating multiple viewpoints [20]. Domain randomization techniques have also been employed to train models that generalize well to real-world scenarios [21]. The common limitation of existing approaches is their unpredictable performance on unseen scenes, highlighting the need for the methodology of automated benchmark data aggregation and annotation without human labor [2].
@@ II-B Existing Datasets
The field of transparent object perception has long faced a scarcity of real-world datasets, with the majority of the work driven by synthetic data [2, 13, 22]. Synthetic datasets like Trans10k [23], SuperCaustics [24], and Dex-NeRF [25] offer a vast number of images (ranging from 9k to 100k) but suffer from the Sim2Real gap [26]. Besides common causes like the lack of noise and imperfections presented in real-world environments, transparent objects provide a unique challenge due to their complex light interaction. In contrast, real-world datasets like TransCG [27] and ClearGrasp [1] have attempted to address these limitations by collecting depth data using physical sensors. However, they lack diversity. Many datasets focus on isolated objects placed against controlled backgrounds, limiting their utility in highly cluttered environments [2]. The process of manual ground-truth depth and segmentation annotation remains a significant bottleneck [28]. Few works, like the Toronto Transparent Object Depth dataset leverage automated annotation techniques [11].
Furthermore, many existing large-scale RGB-D datasets [29] do not specifically focus on transparent objects [30]. We also note the existence of datasets like RGBP-Glass [31] and RGB-Thermal [32], introducing additional sensing modalities. Yet, these remain niche applications and do not fully address the generalization problem for robotic perception.
@@ III-A NICOL Platform and Camera Model
The humanoid robot NICOL, shown in Fig. 1, has two 8-DoF arms that each have an anthropomorphic tendon-driven hand with 5-DoF attached. It has a 2-DoF 3D-printed head structure with an integrated face expression interface. The platform is designed for manipulation and HRI scenarios and offers a variety of sensors for different modalities. A tabletop of two meters in width and one meter in depth is attached in front of NICOL. The robot head contains two RGB fish-eye cameras with a 4k resolution. In our setup, an RGB-D Realsense D435if camera is mounted to NICOL’s head in addition, as well as two more RealSense cameras that are mounted to the left and right front pillars of the robot frame. Thus, the setup enables a complete depth perception of the objects on the table from multiple perspectives.
The moveable head and large area on the table covered by NICOL’s arms, combined with the fish-eye property of the RGB eyes, require a very precise camera calibration to allow for complex manipulation tasks such as pouring liquids. We use a semi-automated calibration method, which scans the table in front of NICOL from different angles to collect calibration data. Small calibration markers are attached to the tabletop that are automatically detected with a blob detection method. The detected positions in the camera image are manually mapped to the corresponding 3D ground truth positions by a human. The collected data is then used with the OpenCV camera calibration method [33] to calculate camera matrix $K$, the distortion coefficients $D$, as well as rotation matrix $R$ and the translation vector $t$. The calibration has to be performed individually for each of the three cameras used in our setup. We reached a reprojection error of $\sim 2~{}pixels$ for the RGB eye cameras and $\sim 4.5~{}pixels$ for the RGB-D camera. The two static RealSense cameras were calibrated with the well-known chessboard calibration method. Our camera model utilizes the camera parameters to provide a ray-casting method between the table and the camera sensor. In other words, pixels from one camera can be cast to the corresponding 3D coordinate at the tabletop and, for example back into one of the other cameras to mark a specific 3D coordinate in the images from all five cameras.
@@ III-B Dataset Design
Recent approaches dealing with transparent object perception usually rely on synthetically rendered data [1, 3, 20]. The amount of real-world data from scenes containing these objects is severely limited, in comparison to datasets with objects from opaque materials [29, 2]. This is especially true for scenes with cluttered environments and a mix of transparent and non-transparent objects. We have designed a new benchmark dataset that is applicable to several computer vision and robotic tasks. Frames are captured from the egocentric view of the NICOL robot [34] performing various arm movements. Thus, complexity in various modalities is added as the dataset includes various degrees of occlusion, varying lighting conditions, and differing tabletop textures.
Specifically, our dataset consists of scenes with a mix of objects placed on a table in front of a humanoid robot. See Fig. 2 for an example scene. The collection process involves three separate passes:
@@
This design allows us to provide glass classification and detection labels, segmentation masks, as well as ground-truth depth measurements. Using the processing pipeline described in the following subsection, all annotations are created automatically, without a need for human labor.
Our setup consists of three RGB-D Realsense cameras, with one attached to the robot’s head and two at the sides of the table. Furthermore, there are two fisheye RGB cameras, one in each eye of the robot, as mentioned earlier. Using this setup, we captured 100 different scenes, with variations in color and texture of the table, lightning conditions, robot movements, and scene compositions. For each scene, there are 25 different views from the robot, acquired by a rotation in the neck joint. All cameras were calibrated and registered, allowing a projection of detection labels from the head-mounted RGB-D sensors to all other views.
In total, the sum of frames acquired from all five camera sensors (2 static RGB-D cameras and 25 images times 3 head cameras) results in a number of 77 frames per scene and 7700 frames in the training dataset. There are six different glass types present, each with its own class, namely: high beer glass, beer glass with a handle, wine glass, water glass, whisky glass, and shot glass. For each type, we also provide a ground-truth 3D model usable in surface reconstruction tasks [35]. Together with ground truth depth observations from the three RGB-D scanners, this can also be used in monocular depth estimation [36] and depth completion tasks [27]. Our main use case of the data presented in this paper is the detection and classification of a glass type. In a more general perception task, different glass types can be omitted, and all labels can be treated as general drinking glass annotations. Our data intentionally contained various unlabeled bottles. These provide negative examples in training for objects that are made from transparent material but would not be classified as any type of glass for drinking by a human.
In addition to the 7700 frames used for training and validation, we have also captured separate testing data. It contains 150 frames, with 75 frames captured using the same approach as the training data, plus another 75 frames captured during the bartender experiment described in Sec. IV-B, thus having a completely different character compared to the training data. Labels for the test data have been created manually, to mitigate the effect of label noise introduced by our automated pipeline described below.
@@ III-C Auto-Labeling Pipeline
$h=\frac{|d^{\prime}-d|}{\sqrt{a^{2}+b^{2}+c^{2}}}$ (1)
where $d^{\prime}$ is the cluster’s plane offset. We discard clusters with large height deviations from known glasses. The remaining candidates are projected onto the 2D image plane using intrinsic camera parameters, transforming 3D coordinates into pixel locations $(u,v)$ through perspective projection:
$\begin{bmatrix}u&v&1\end{bmatrix}^{T}\sim K\begin{bmatrix}x&y&z&1\end{bmatrix}^{T}$ (2)
where $K$ is the intrinsic calibration matrix of the camera. To further filter instances, color verification is performed in the CIELAB color space, ensuring that the detected object caps match the expected green hue. A final verification step utilizes YOLO-World, a deep-learning-based object detector, to confirm the detected object has glass-like features, mitigating false positives from other green objects in the scene [7]. As each of these steps work with different data modalities and different false-positive error profiles, we achieve a variant of a cascaded filtering approach by sensor fusion [38]. After obtaining the final candidates, points from the blobs are used as samples for the Segment Anything Model (SAM) [39], generating object masks for precise annotations. The segmentation masks $M_{i}$ are converted into polygons, defining the object’s shape in the image space. We derive bounding boxes $B_{i}=(x,y,w,h)$ from these masks, finalizing the object instance label. A visual overview of our labeling pipeline and an example annotation were presented earlier (see Fig. LABEL:fig:pipe).
@@ III-D Baseline Detector
To demonstrate the advantages of our novel dataset, we use it to train an object detector and perform use-case scenarios and experiments. For real-time detection performance, we opted for an RTMDet architecture in its medium-size variant [40]. Naturally, different one-stage architectures, like the YOLO family [41], could be employed for this task.
The chosen network variant incorporates various training strategies. Specifically, the network employs a CSPNeXt backbone with a P5 architecture, a deepen factor of 0.67, and a widen factor of 0.75, enhanced with channel attention and synchronized batch normalization. The neck is a CSPNeXtPAFPN structure with two CSP blocks [42], facilitating multi-scale feature aggregation [40]. As proposed by the original authors, the head of the network calculates three loss functions. A Quality Focal Loss ($\beta=1.0$, loss weight = 1.0) for classification [43], Complete IoU (CIoU) Loss (loss weight = 2.0) for bounding box regression [44], and Dice Loss (loss weight = 3.0, reduction = “mean”) for mask prediction [40]. To increase robustness and grant better generalization for different scenarios, heavy data augmentation is employed. This includes techniques such as Cached Mosaic [45], Cached MixUp [46], Random Resize, Random Crop, and YOLOX-style HSV augmentation [47].
The training is performed from a pre-trained checkpoint [40], hence we only fine-tune it on our data. The learning rate is scheduled using a combination of Linear Warmup (start factor = 0.001, epochs 0–50) and MultiStepLR with decay at epochs 100, 200, and 400. The optimizer is AdamW [48] with a learning rate of 0.01 and a weight decay of 0.05. The model is trained for a total of 500 epochs with a batch size of 8, and evaluation is conducted every 50 epochs using the COCO metric for both bounding box and segmentation performance on 5 (representing a 5% split from our 100 scenes) validations scenes.

### v3

@@ Abstract
Datasets for object detection often do not account for enough variety of glasses, due to their transparent and reflective properties. Specifically, open-vocabulary object detectors, widely used in embodied robotic agents, fail to distinguish subclasses of glasses. This scientific gap poses an issue for robotic applications that suffer from accumulating errors between detection, planning, and action execution. This paper introduces a novel method for acquiring real-world data from RGB-D sensors that minimizes human effort. We propose an auto-labeling pipeline that generates labels for all the acquired frames based on the depth measurements. We provide a novel real-world glass object dataset [^3] https://knowledgetechnologyuhh.github.io/GlassNICOLDataset/ that was collected on the Neuro-Inspired COLlaborator (NICOL), a humanoid robot platform. The dataset consists of 7850 images recorded from five different cameras. We show that our trained baseline model outperforms state-of-the-art open-vocabulary approaches. In addition, we deploy our baseline model in an embodied agent approach to the NICOL platform, on which it achieves a success rate of 81% in a human-robot bartending scenario.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I INTRODUCTION
Transparent objects are everywhere, from households, healthcare assistance, and gastronomy to industrial and construction-site environments. Yet, these materials reflect only a very small fraction of visible light—most of it is scattered and passes through. From a computer vision perspective, the processing of these objects is challenging but crucial for a successful deployment in real-world robotics applications [1]. As intelligent robots start to make their way into public spaces, where glass objects in the form of bottles and drinking glasses are integral items, future generations of service robots will be required to not only detect but also manipulate those materials. In addition, machine learning and data-driven methods require a large amount of training data to reach a suitable performance. Similar to robotic Sim2Real applications, existing research also proves the presence of a significant Sim2Real gap in the area of synthetic glass material generation [2, 3, 1], while those approaches require sophisticated photorealistic rendering methods.
Open‑vocabulary approaches are used for object detection in a plethora of embodied robotic setups, such as [4, 5]. Embodied robotic agents, at a minimum, consist of a detection module, an action module, and an LLM utilized for high-level interaction and task planning. Recent approaches aim to fuse detection and language processing by utilizing vision-language models (VLMs) [6]. Being interactive by design, these robotic agents are typically situated in scenarios involving user interaction and often collaboration, leaving a potentially infinite corpus of object names to be detected.
Well-known open-vocabulary object detection approaches such as Grounding DINO [7], YOLOWorld [8], and OWL-Vit [9] have revolutionized object detection by expanding the recognizable object classes without extensive re‑annotation. These methods are built upon end-to-end transformer‑based architectures. Global image features are extracted by decoder backbones and fused with language embeddings by utilizing deep metric learning. Despite their success, our experiments suggest that these models exhibit a significant decrease in performance when encountering transparent objects, commonly found in HRI scenarios. To address these challenges, we introduce a novel real-world dataset that was captured on the humanoid robot NICOL (Fig. 1), including a new depth-based automated labeling method. The robot is equipped with a combination of RGB-D scanners and standard RGB cameras. Specific contributions of our work include:
@@ II-A Transparent Material Perception
Missing or incorrect RGB-D data for robotic perception can be estimated using implicit functions and geometric priors [10] and by balancing local and global depth features [11]. The depth map can also be de-projected into a point cloud and processed with techniques like 3D CNNs to achieve depth completion [12]. Iterative refinement is also possible, with an attention module on transparent regions [13] or with indirect geometry representations [14].
RGB often contains information missing from depth data, which can be leveraged with transfer learning [15]. Cues can be extracted from the RGB to fill in depth observations using global optimization [1]. Incorporating affordance detection of the underlying object yields further improvement [16]. Other approaches include multimodality, such as visual-tactile fusion [17] and usage of polarization imaging [18].
Recently, neural radiance fields (NeRFs), have emerged as a promising tool for transparent object perception. GraspNeRF introduces a multiview-based 6-DoF grasp detection system [19]. Evo-NeRF extends this approach with real-time NeRF training and grasp adaptation [20]. Multiview methods enhance transparent object perception by integrating multiple viewpoints [21]. Domain randomization techniques have also been used to train models that generalize well to the real-world [22]. The common limitation of existing approaches is their unpredictable performance on unseen scenes, showing the need for the methodology of automated benchmark data aggregation and annotation without human labor [2].
@@ II-B Existing Datasets
The field of transparent object perception has long faced a scarcity of real-world datasets, with the majority of the work driven by synthetic data [2, 14, 23]. Synthetic datasets like Trans10k [24], SuperCaustics [25], and Dex-NeRF [26] offer a vast number of images (ranging from 9k to 100k) but suffer from the Sim2Real gap [27]. Besides common causes like the lack of noise and imperfections presented in real-world environments, transparent objects provide a unique challenge due to their complex light interaction. In contrast, real-world datasets like TransCG [28] and ClearGrasp [1] have attempted to address these limitations by collecting depth data using physical sensors. However, they lack diversity. Many datasets focus on isolated objects placed against controlled backgrounds, limiting their utility in highly cluttered environments [2]. The process of manual ground-truth depth and segmentation annotation remains a significant bottleneck [29]. Few works, like the Toronto Transparent Object Depth dataset leverage automated annotation techniques [12]. Also, Liu et al. [30] presented a multiview automated annotation approach. However, it is limited to isolated scenes and does not address the demanding challenge of dense transparent object detection and classification in realistic HRI scenarios, which our dataset is specifically designed for.
Moreover, many existing large-scale RGB-D datasets [31] do not specifically focus on transparent objects [32]. We also note the existence of datasets like RGBP-Glass [33] and RGB-Thermal [34], introducing additional sensing modalities. Yet, these remain niche applications and do not fully address the generalization problem for robotic perception.
@@ III-A NICOL Platform and Camera Model
The humanoid robot NICOL (see Fig. 1) has two 8-DoF arms, each with an anthropomorphic tendon-driven 5-DoF hand. It has a 2-DoF 3D-printed head structure with an integrated facial expression interface. The platform is designed for manipulation and HRI scenarios and offers various sensors for different modalities. A two-meter-wide, one-meter-deep tabletop is attached in front of NICOL. The robot head contains two RGB fish-eye cameras with a 4k resolution. In our setup, a RealSense D435if camera is additionally mounted to NICOL’s head, along with two RealSense cameras to the left and right front pillars of the robot’s frame. Thus, our setup enables a complete depth perception of the objects on the table from multiple perspectives.
The movable head and large area on the table covered by NICOL’s arms, combined with the fish-eye property of the RGB eyes, require a very precise camera calibration to allow for complex manipulation tasks such as pouring liquids. We use a semi-automated calibration method, which scans the table in front of NICOL from different angles to collect calibration data. Small calibration markers are attached to the tabletop that are automatically detected with a blob detection method. The detected positions in the camera image are manually mapped to the corresponding 3D ground truth positions by a human. The collected data is then used with the OpenCV camera calibration method [35] to calculate the camera matrix $K$, the distortion coefficients $D$, as well as the rotation matrix $R$ and the translation vector $t$. The calibration has to be performed individually for each of the three cameras used in our setup. We reached a reprojection error of $\sim 2~pixels$ for the RGB eye cameras and $\sim 4.5~pixels$ for the RGB-D camera. The two static RealSense cameras were calibrated with the well-known chessboard calibration method. Our camera model utilizes the camera parameters to provide a ray-casting method between the table and the camera sensor. In other words, pixels from one camera can be cast to the corresponding 3D coordinate at the tabletop and, for example, back into one of the other cameras to mark a specific 3D coordinate in the images from all five cameras.
@@ III-B Dataset Design
Recent approaches dealing with transparent object perception usually rely on synthetically rendered data [1, 3, 21]. The amount of real-world data from scenes containing these objects is severely limited, in comparison to datasets with objects from opaque materials [31, 2]. This is especially true for scenes with cluttered environments and a mix of transparent and non-transparent objects. We have designed a new benchmark dataset that is applicable to several computer vision and robotic tasks. Image frames are captured from the egocentric view of the NICOL robot [36] performing various arm movements. Thus, complexity in various modalities is added as the dataset includes various degrees of occlusion, varying lighting conditions, and differing tabletop textures.
Specifically, our dataset consists of scenes with a mix of objects placed on a table in front of a humanoid robot (see Fig. 2). The collection process involves three separate passes:
@@
This design allows us to provide glass classification and detection labels, segmentation masks, and ground-truth depth measurements. Using the processing pipeline described in the following subsection, all annotations are created automatically, without human labor. As mentioned earlier, our setup consists of three RealSense RGB-D cameras, one attached to the robot’s head and two at the sides of the table. Also, there are two fisheye RGB cameras, one in each eye of the robot. Using this setup, we captured 100 different scenes, with variations in table color and texture, lightning conditions, robot movements, and scene compositions. For each scene, there are 25 different views from the robot, acquired by a rotation in the neck joint. All cameras were calibrated and registered, allowing a projection of detection labels from the head-mounted RGB-D sensors to all other views.
The sum of frames acquired from all five cameras (2 static RGB-D cameras and 25 images times 3 head cameras) results in a number of 77 frames per scene and 7700 frames in the training dataset in total. There are six different glass types present, each with its own class, namely: high beer glass, beer glass with a handle, wine glass, water glass, whiskey glass, and shot glass. For each type, we also provide a high-quality 3D model usable in surface reconstruction tasks [37]. Together with ground truth depth observations from the three RGB-D scanners, this can also be used in monocular depth estimation [38] and depth completion tasks [28]. Our main use case of the data in this paper is the detection and classification of a glass type. In a more general perception task, different glass types can be omitted, and all labels can be treated as general drinking glass annotations.
Our dataset intentionally contains various unlabeled bottles, serving as negative examples during training. These objects are made of transparent material but are not typically classified by humans as drinking glasses. In addition to the 7700 frames used for training and validation, we captured separate testing data containing 150 frames: 75 frames captured using the same approach as the training data and another 75 frames captured during the bartender experiment described in Sec. IV-B, thus having a completely different character compared to the training data. The labels of the test data were created manually to mitigate the effect of label noise introduced by our automated pipeline described below.
@@ III-C Auto-Labeling Pipeline
$h=\frac{|d^{\prime}-d|}{\sqrt{a^{2}+b^{2}+c^{2}}},$ (1)
where $d^{\prime}$ is the cluster’s plane offset. Clusters with large height deviations from known glasses are discarded. The remaining candidates are given classes based on the closest known glass height, then projected onto the 2D image plane using intrinsic camera parameters, transforming 3D coordinates into pixel locations $(u,v)$ via perspective projection:
$\begin{bmatrix}u&v&1\end{bmatrix}^{T}\sim K\begin{bmatrix}x&y&z&1\end{bmatrix}^{T},$ (2)
where $K$ is the intrinsic calibration matrix of the camera. To further filter instances, color verification is performed in the CIELAB color space, ensuring that the detected object caps match the expected green hue. A final verification step uses YOLO-World, a deep-learning-based object detector, to confirm the detected object has glass-like features, mitigating false positives from other green objects in the scene [8]. As each of these steps works with different data modalities and false-positive error profiles, we achieve a variant of a cascaded filtering approach by fusion of outputs from several algorithms  [40]. After obtaining the final candidates, points from the blobs are used as samples for the Segment Anything Model (SAM) [41]. As shown by our final data, SAM achieves reliable segmentation performance in cluttered environments when guided by our precise sample points. To improve accuracy, we discard overly wide masks, often caused by merged neighboring objects. To finalize the object instance label, we derive bounding boxes $B_{i}=(x,y,w,h)$ from the masks. An overview of our labeling pipeline and an example annotation were presented earlier (see Fig. LABEL:fig:pipe).
@@ III-D Baseline Detector
To demonstrate the advantages of our novel dataset, we use it to train an object detector and perform use case scenarios and experiments. For real-time detection performance, we opted for an RTMDet architecture in its medium-size variant [42]. Naturally, different one-stage architectures, like the YOLO family [43], could be employed for this task.
The network variant chosen incorporates various training strategies. Specifically, the network employs a CSPNeXt backbone with a P5 architecture, a deepen factor of 0.67, and a widen factor of 0.75, enhanced with channel attention and synchronized batch normalization. The neck is a CSPNeXtPAFPN structure with two CSP blocks [44], facilitating multi-scale feature aggregation [42]. As proposed by the original authors, the head of the network calculates three loss functions. A Quality Focal Loss ($\beta=1.0$, loss weight = 1.0) for classification [45], Complete IoU (CIoU) Loss (loss weight = 2.0) for bounding box regression [46], and Dice Loss (loss weight = 3.0, reduction = “mean”) for mask prediction [42]. To increase robustness and grant better generalization for different scenarios, heavy data augmentation is employed. This includes techniques such as Cached Mosaic [47], Cached MixUp [48], Random Resize, Random Crop, and YOLOX-style HSV augmentation [49].
The training is performed from a pre-trained checkpoint [42], which we fine-tune on our dataset. The learning rate is scheduled using a combination of Linear Warmup (start factor = 0.001, epochs 0–50) and MultiStepLR with decay at epochs 100, 200, and 400. The optimizer is AdamW [50] with a learning rate of 0.01 and a weight decay of 0.05. The model is trained for 500 epochs with a batch size of 8, and evaluated every 50 epochs using the COCO metric for both bounding box and segmentation performance on 5 validation scenes (a 5% split from our 100 scenes).